# Demystifying amortized causal discovery with transformers

### v1

@@ 3.2 Warm up: is CSIvA capable of in and out-of-distribution generalization?
First, we investigate the generalization of CSIvA on datasets sampled from the structural casual model that generates the train distribution, with mechanisms and noise distributions fixed between training and testing. We call this in-distribution generalization. As a benchmark, we present the performance of several state-of-the-art approaches from the literature on causal discovery: we consider the DirectLiNGAM, and NoGAM algorithms [34, 35], respectively designed for the inference on LiNGAM and nonlinear ANM generated data [^1] The causal-learn implementation of the PNL algorithm could not perform better than random on our synthetic post-nonlinear data, and we observed that this was due to the sensitivity of the algorithm to the variance scale. So we report the plot of Figure 1(c) without benchmark comparison. We remark that the point of this experiment is not to make any claims on CSIvA being state-of-the-art but to validate that the performance we obtain in our re-implementation is non-trivial. This is clear for PNL, even without comparison.. The results of Figure 1 show that CSIvA can properly generalize to unseen samples from the training distribution: the majority of the trained models present SHD close to zero and comparable to the relative benchmark algorithm.
@@
Figure 1: In-distribution generalization of CSIvA trained and tested on data generated according to the same structural causal models, fixing mechanisms, and noise distributions between training and testing). As baselines for comparison, we use DirectLiNGAM on linear SCMs and NoGAM on nonlinear ANM (we use their causal-learn and dodiscover implementations). CSIvA performance is clearly non-trivial and generalizing well.
@@
In practice, we generally do not know the SCM defining the test distribution, so we are interested in CSIvA’s ability to generalize to data sampled from a class of causal models that is unobserved during training. We call this out-of-distribution generalization (OOD). We study OOD generalization to different noise terms, analyzing the network performance on datasets generated from causal models where the mechanisms are fixed with respect to the training, while the noise distribution varies (e.g., given linear-mlp training samples, testing occurs on linear-uniform data). Orthogonally to these experiments, we empirically validate CSIvA’s OOD generalization over different mechanism types (linear, nonlinear, post-nonlinear), while leaving the noise distribution (mlp) fixed across test and training. In Figure 2(a), we observe that CSIvA cannot generalize across the different mechanisms, as the SHD of a network tested on unseen causal mechanisms approximates that of the random baseline. Further, Figure 2(b) shows that out-of-distribution generalization across noise terms does not work reliably, and it is hard to predict when it might occur.
@@
CSIvA generalizes well to test data generated by the same class of SCMs used for training, in line with the findings in Ke et al. [1], which validates our implementation and training procedure. However, it struggles when the test data are out-of-distribution, not generated by causal models with the same mechanisms and noise terms it was trained on. While training on a wider class of SCMs might overcome this limitation, it requires caution. The identifiability of causal graphs indeed results from the interplay between the data-generating mechanisms and noise distribution. However, as we argue in our Example 1, the class of causal models that a supervised learning algorithm can identify is generally not clear. In what follows, we investigate this point and its implications for CSIvA, showing that the identifiability of the test samples can be ensured by imposing suitable assumptions on the class of SCMs generating the training distribution.
@@
Figure 2: Out-of-distribution generalisation. We train three CSIvA models on data sampled from SCMs with linear, nonlinear additive, and post-nonlinear mechanisms; and noise fixed mlp noise distribution. In Figure (a) we test across different noise distributions, with test mechanisms fixed from training. In Figure (b) we test each network on different mechanisms and fixed mlp noise. CSIvA struggles to generalize to unseen causal mechanisms and often displays degraded performance over new noise distributions.
@@ 3.3 How does CSIvA relate to identifiability theory for causal graphs?
The CSIvA algorithm does not make structural assumptions about the causal model underlying the input data. This implies that the output of this method is unclear: as CSIvA targets the conditional distribution $p(\cdot|\mathcal{D})$ over the space of graphs, in the absence of restrictions on the functional mechanisms and the distribution of the noise terms, the causal graph $X\to Y$ is indistinguishable from $Y\to X$, as they are both equally likely to underlie the joint density $p_{X,Y}$ generating the data. As we discuss in Example 1, the graphical output of the trained architecture could at most identify the equivalence class of the true causal graph. Yet, our experiments of Section 3.2 show that CSIvA is capable of good in-distribution generalization, often inferring the correct DAG at test time. We explain this seeming contradiction with the following hypothesis, which motivates the analysis in the remainder of this section.
@@
To support and clarify our statement, we present the following example, adapted from Hoyer et al. [7].
@@
Consider the causal model $Y=f(X)+N,$ where $f(X)=-X$ and $p_{X},p_{N}$ are Gumbel densities $p_{X}(x)=\operatorname{exp}(-x-\operatorname{exp}(-x))$ and $p_{N}(n)=\operatorname{exp}(-n-\operatorname{exp}(-n))$. This model satisfies the assumptions of the LiNGAM, so it is identifiable, in the sense that a backward linear model with the same distribution does not exist. However, in this special case, we can build a backward nonlinear additive noise model $X=g(Y)+\tilde{N}$ with independent noise terms: taking $p_{Y}(y)=\operatorname{exp}(-y-2\log(1+\operatorname{exp}(-y)))$ to be the density of a logistic distribution, $p_{\tilde{N}}(\tilde{n})=\operatorname{exp}(-2\tilde{n}-\operatorname{exp}(-\tilde{n}))$ and $g(y)=\log(1+\operatorname{exp}(-y))$; we see that $p_{X,Y}$ can factorize according to two opposite causal directions, as $p_{X,Y}(x,y)=p_{N}(y-f(x))p_{X}(x)=p_{\tilde{N}}(x-g(y))p_{Y}(y)$. Given a dataset $\mathcal{D}$ of observations from the forward linear model, causal discovery methods like DirectLiNGAM [34] can provably identify the correct causal direction $X\rightarrow Y$, assuming that sufficient samples are provided. Instead, the behavior of CSIvA seems hard to predict: given that the network approximates the conditional distribution $p(\cdot|\mathcal{D})$ over the possible graphs, for $\mathcal{D}$ with arbitrary many samples we have $p(X\rightarrow Y|\mathcal{D})=p(Y\rightarrow X|\mathcal{D})=0.5$. On the other hand, given the prior knowledge that the data-generating SCM is a linear non-gaussian additive noise model, we have $p(X\rightarrow Y|\mathcal{D},\textnormal{LiNGAM})=1$, because the LiNGAM is identifiable. In this sense, the class of structural causal models that CSIvA correctly infers appears to be determined by the structural causal models underlying the generation of the training data. Under our Hypothesis 1, training CSIvA exclusively on LiNGAM-generated data is equivalent to learning the distribution $p(\cdot|\mathcal{D},\textnormal{LiNGAM})$, such that the network should be able to identify the forward linear model, whereas it could only infer the equivalence class of the causal graph if its training datasets include observations from a nonlinear additive noise model.
@@
Figure 3: Experiments on identifiability theory. In Figure (a) we test the performance on linear-Gaussian data. Models are trained with different ratios of samples from linear and nonlinear SCMs with Gaussian noise terms. The validation results showcase that the networks were trained successfully. Figure (b) shows the SHD of models trained on different ratios of linear and nonlinear invertible data of Example 2. CSIvA behaves according to identifiability theory, failing to predict on linear Gaussian models and invertible data (50:50 ratio).
@@
Our experiments show that CSIvA learns algorithms that closely follow identifiability theory for causal discovery. In particular, while the method itself does not require explicit assumptions on the data-generating process, the chosen training data ultimately determines the class of causal models identifiable during inference. Notably, previous work has argued that supervised learning approaches in causal discovery would help with "dealing with complex data-generating processes and greatly reduce the need of explicitly crafting identifiability conditions a-priori", Lopez-Paz et al. [10]. In the case of CSIvA, this expectation does not appear to be fulfilled, as the assumptions still need to be encoded explicitly in the training data. However, this observation opens two new important questions: (1) Can we train a single network to encompass multiple (or even all) identifiable causal structures? (2) How much ambiguity might exist between these identifiable models?
3.4 A low-dimensions argument in favor of learning from multiple causal models
Example 2 of the previous section shows that elements of distinct classes of identifiable structural causal models, such as LiNGAM and nonlinear ANM, may become non-identifiable when we consider their union. In this section, we show that in the class of post-additive noise models given by equation (2) (obtained as the union of the LiNGAM, the nonlinear ANM, and the post-nonlinear model), the set of distributions that is non-identifiable is negligible. Our proposition extends the results of Hoyer et al. [7], which are limited to the case of linear and nonlinear additive noise models, and Zhang and Hyvärinen [8], which provides the conditions of identifiability of the post-ANM without bounding the set of non-identifiable distributions.
@@ 3.4 A low-dimensions argument in favor of learning from multiple causal models
We note that this implies that the following invertible additive noise models on $\tilde{X},\tilde{Y}$ hold:
@@
Proposition 1 (Adapted from Hoyer et al. [7]).
@@
Our result is closely related to Theorem 1 of Hoyer et al. [7], which we simply generalize to the post-ANM. Intuitively, it says that the space of all continuous distributions such that the bivariate post-ANM is non-identifiable is contained in a 2-dimensional space. As the space of continuous distributions of random variables is infinite-dimensional, we conclude that the post-ANM is generally identifiable, which suggests that the setting of Example 2 is rather artificial. Our results provide a theoretical ground for training causal discovery algorithms on datasets generated from multiple identifiable SCMs. This is particularly appealing in the case of CSIvA, given the poor OOD generalization ability observed in our experiments of Section 3.2.
@@ 3.5 Can we train CSIvA on multiple causal models for better generalization?
Next, we analyze the test performance of three CSIvA networks optimized on samples from structural causal models that have different distributions for their noise terms, while keeping the mechanism types fixed. Figure 5 shows that training over different noises (beta, gamma, gumbel, exponential, mlp, uniform) always results in a network that is agnostic with respect to the noise distributions of the SCM generating the test samples, always achieving SHD $<0.1$, with the exception of datasets with mlp error terms ($0.2$ average SHD on nonlinear and pnl data).
@@
We have shown that learning on mixtures of SCMs with different noise term distributions and mechanism types leads to models generalizing to a much broader class of structural causal models during testing. Hence, combining datasets generated from multiple models looks like a promising framework to overcome the limited out-of-distribution generalization abilities of CSIvA observed in Section 3.2. However, it is easier to incorporate prior assumptions on the class of causal mechanisms (linear, non-linear, post-non-linear) compared to the noise distributions (which are potentially infinite). This introduces a trade-off between amortized inference and classical methods for causal discovery: for example, RESIT, NoGAM, and CAM [23, 35, 36] algorithms require no assumptions on the noise type, but only work for a limited class of mechanisms (nonlinear).
@@ 4 Conclusion
In this work, we investigate the interplay between identifiability theory and supervised learning for amortized inference of causal graphs, using CSIvA as the ground of our study. Consistent with classical algorithms, we demonstrate that good performance can be achieved if (i) we have a good prior on the structural causal model generating the test data (ii) the setting is identifiable. In particular, prior knowledge of the test distribution is encoded in the training data in the form of constraints on the structural causal model underlying their generation. With these results, we highlight the need for identifiability theory in modern learning-based approaches to causality, while past works have mostly disregarded this connection. Further, our findings provide the theoretical ground for training on observations sampled from multiple classes of identifiable SCMs, a strategy that improves test generalization to a broad class of causal models. Finally, we highlight an interesting new trade-off regarding identifiability: traditional methods like LiNGAM, RESIT, and PNL require strong restrictions on the structural mechanisms underlying the data generation (linear, nonlinear or post-nonlinear) while generally being agnostic relative to the noise terms distribution. Training on mixtures of causal models instead offers an alternative that is less reliant on assumptions on the mechanisms, while incorporating knowledge about all possible noise distributions in the training data is practically impossible to achieve. We leave it to future work to reproduce our analysis on a wider class of architectures, as well as extending our study to interventional data with more than two nodes.

### v2

@@ 4.2 Warm up: is CSIvA capable of in and out-of-distribution generalization?
First, we investigate the generalization of CSIvA on datasets sampled from the structural casual model that generates the train distribution, with mechanisms and noise distributions fixed between training and testing. We call this in-distribution generalization. The main goal of these experiments is to validate that the performance of our CSIvA implementation is non-trivial. As a benchmark, we present the accuracy of two state-of-the-art approaches from the literature on causal discovery: we consider the DirectLiNGAM and NoGAM algorithms (Shimizu et al., 2011, Montagna et al., 2023c), respectively designed for the inference on LiNGAM and nonlinear ANM generated data [^1] The causal-learn implementation of the PNL algorithm could not perform better than random on our synthetic post-nonlinear data, and we observed that this was due to the sensitivity of the algorithm to the variance scale. So we report the plot of Figure 1(c) without benchmark comparison. We remark that the point of this experiment is not to make any claims on CSIvA being state-of-the-art but to validate that the performance we obtain in our re-implementation is non-trivial. This is clear for PNL, even without comparison.. The results of Figure 1 show that CSIvA can properly generalize to unseen samples from the training distribution: the majority of the trained models present SHD close to zero and comparable to the relative benchmark algorithm.
@@
Figure 1: In-distribution generalization of CSIvA trained and tested on data generated according to the same structural causal models, fixing mechanisms, and noise distributions between training and testing. As baselines for comparison, we use DirectLiNGAM on linear SCMs and NoGAM on nonlinear ANM (we use their causal-learn and dodiscover implementations). CSIvA performance is clearly non-trivial and generalizing well.
@@
In practice, we generally do not know the SCM defining the test distribution, so we are interested in CSIvA’s ability to generalize to data sampled from a class of causal models that is unobserved during training. We call this out-of-distribution generalization (OOD). We study OOD generalization to different noise terms, analyzing the network performance on datasets generated from causal models where the mechanisms are fixed with respect to the training, while the noise distribution varies (e.g., given linear-mlp training samples, testing occurs on linear-uniform data). Orthogonally to these experiments, we empirically validate CSIvA’s OOD generalization over different mechanism types (linear, nonlinear, post-nonlinear), while leaving the noise distribution (mlp) fixed across test and training. In Figure 2(a), we observe that CSIvA cannot generalize across the different mechanisms, as the SHD of a network tested on unseen causal mechanisms approximates that of the random baseline. Further, Figure 2(b) shows that out-of-distribution generalization across noise terms does not work reliably, and it is hard to predict when it might occur. We note that these findings are novel in the literature: OOD experiments in the sense we define can not be found in Ke et al. (2023b) in the first place; Li et al. (2020) empirical results are limited to the OOD generalization from linear Gaussian models to linear models with Exponential, Gumbel and Poisson noise. Lorch et al. (2022) analyses generalization from SCMs with Gaussian noise to Laplace and Cauchy distributions, and fixed mechanisms class. The other works on amortized inference in our related works section 2 generally disregard these experiments.
@@
CSIvA generalizes well to test data generated by the same class of SCMs used for training, in line with the findings in Ke et al. (2023b), which validates our implementation and training procedure. However, it struggles when the test data are out-of-distribution, generated by causal models with different mechanism types and noise distributions that those it was trained on. From a practical perspective, this is a relevant finding, given that existing works on amortized causal discovery lacks systematic experiments on the OOD setting. While training on a wider class of SCMs might overcome this limitation, it requires caution. The identifiability of causal graphs indeed results from the interplay between the data-generating mechanisms and noise distribution. However, as we argue in our Example 1, the class of causal models that a supervised learning algorithm can identify is generally not clear. In what follows, we investigate this point and its implications for CSIvA, showing that the identifiability of the test samples can be ensured by imposing suitable assumptions on the class of SCMs generating the training distribution.
@@
Figure 2: Out-of-distribution generalisation. We train three CSIvA models on data sampled from SCMs with linear, nonlinear additive, and post-nonlinear mechanisms; and fixed mlp noise distribution. In Figure 2(a) we test across different mechanism types, with mlp-distributed noise terms both in test and training. In Figure 2(b) we test across different noise distributions, with test mechanism types fixed from training. CSIvA struggles to generalize to unseen causal mechanisms and often displays degraded performance over new noise distributions.
@@ 4.3 How does CSIvA relate to identifiability theory for causal graphs?
The CSIvA algorithm does not make structural assumptions about the causal model underlying the input data. This implies that the output of this method is unclear: as CSIvA should target the conditional distribution $p(\cdot|\mathcal{D})$ over the space of graphs, in the absence of restrictions on the functional mechanisms and the distribution of the noise terms, the causal graph $X\to Y$ is indistinguishable from $Y\to X$, as they are both equally likely to underlie the joint density $p_{X,Y}$ generating the data. As we discuss in Example 1, the graphical output of the trained architecture could at most identify the equivalence class of the true causal graph. Yet, our experiments of Section 4.2 show that CSIvA is capable of good in-distribution generalization, often inferring the correct DAG at test time. We explain this seeming contradiction with the following hypothesis, which motivates the analysis in the remainder of this section.
@@
To support and clarify our statement, we present the following example, adapted from Hoyer et al. (2008).
@@
Consider the causal model $Y=f(X)+N,$ where $f(X)=-X$ and $p_{X},p_{N}$ are Gumbel densities $p_{X}(x)=\operatorname{exp}(-x-\operatorname{exp}(-x))$ and $p_{N}(n)=\operatorname{exp}(-n-\operatorname{exp}(-n))$. This model satisfies the assumptions of the LiNGAM, so it is identifiable, in the sense that a backward linear model with the same distribution does not exist. However, in this special case, we can build a backward nonlinear additive noise model $X=g(Y)+\tilde{N}$ with independent noise terms: taking $p_{Y}(y)=\operatorname{exp}(-y-2\log(1+\operatorname{exp}(-y)))$ to be the density of a logistic distribution, $p_{\tilde{N}}(\tilde{n})=\operatorname{exp}(-2\tilde{n}-\operatorname{exp}(-\tilde{n}))$ and $g(y)=\log(1+\operatorname{exp}(-y))$; we see that $p_{X,Y}$ can factorize according to two opposite causal directions, as $p_{X,Y}(x,y)=p_{N}(y-f(x))p_{X}(x)=p_{\tilde{N}}(x-g(y))p_{Y}(y)$. Given a dataset $\mathcal{D}$ of observations from the forward linear model, causal discovery methods like DirectLiNGAM (Shimizu et al., 2011) can provably identify the correct causal direction $X\rightarrow Y$, assuming that sufficient samples are provided. Instead, the behavior of CSIvA seems hard to predict: given that the network approximates the conditional distribution $p(\cdot|\mathcal{D})$ over the possible graphs, for $\mathcal{D}$ with arbitrary many samples we have $p(X\rightarrow Y|\mathcal{D})=p(Y\rightarrow X|\mathcal{D})=0.5$. On the other hand, given the prior knowledge that the data-generating SCM is a linear non-gaussian additive noise model, we have $p(X\rightarrow Y|\mathcal{D},\textnormal{LiNGAM})=1$, because the LiNGAM is identifiable. In this sense, the class of structural causal models that CSIvA correctly infers appears to be determined by the structural causal models underlying the generation of the training data. Under our Hypothesis 1, training CSIvA exclusively on LiNGAM-generated data is equivalent to learning the distribution $p(\cdot|\mathcal{D},\textnormal{LiNGAM})$, such that the network should be able to identify the forward linear model, whereas it could only infer the equivalence class of the causal graph if its training datasets include observations from a nonlinear additive noise model.
@@
Figure 3: Experiments on identifiability theory. Figure 3(a) shows the SHD of models trained on different ratios of linear and nonlinear invertible data of Example 2. In Figure 3(b) we test the performance on linear-Gaussian data. Models are trained with different ratios of samples from linear and nonlinear SCMs with Gaussian noise terms. The validation results showcase that the networks were trained successfully. In both cases CSIvA behaves according to identifiability theory, failing to predict on invertible data (50:50 ratio) and linear Gaussian models.
@@
Our experiments show that CSIvA learns algorithms that closely follow identifiability theory for causal discovery. In particular, while the method itself does not require explicit assumptions on the data-generating process, the chosen training data ultimately determines the class of causal models identifiable during inference. Notably, previous work has argued that supervised learning approaches in causal discovery would help with "dealing with complex data-generating processes and greatly reduce the need of explicitly crafting identifiability conditions a-priori", Lopez-Paz et al. (2015). In the case of CSIvA, this expectation does not appear to be fulfilled, as the assumptions still need to be encoded explicitly in the training data. However, this observation opens two new important questions: (1) Can we train a single network to encompass multiple (or even all) identifiable causal structures? (2) How much ambiguity might exist between these identifiable models? We start by answering this second question.
4.4 An identifiability argument in favor of learning from multiple causal models
Example 2 of the previous section shows that elements of distinct classes of identifiable structural causal models, such as LiNGAM and nonlinear ANM, may become non-identifiable when we consider their union. In this section, we discuss the identifiability of the post-additive noise models. Previously, Hoyer et al. (2008) showed that the set of distributions generated according to the additive noise model (3) and that is non-identifiable is negligible. Later, Zhang & Hyvärinen (2009) characterized non-identifiable post-nonlinear models in terms of the properties of their functional mechanisms, and the distribution of the noise terms. In this section, we discuss how these results put together show that the set of distributions generated according to a post-ANM that is non-identifiable is negligible.
@@ 4.4 An identifiability argument in favor of learning from multiple causal models
We note that equivalently the following invertible additive noise models on $\tilde{X},\tilde{Y}$ hold:
@@
Equations (7) and (8) reduce the problem of studying the identifiability of a post-ANM to that of studying the identifiability of an additive noise model, as done in Theorem 1 of Hoyer et al. (2008), which we repropose in Appendix E: intuitively, the statement of the theorem says that the space of all continuous distributions generated according to a bivariate additive noise model and that is non-identifiable is contained in a 2-dimensional space. As the space of continuous distributions of random variables is infinite-dimensional, we conclude that the ANM is generally identifiable. Given that, according to Equation (7) and Equation (8), the post-ANM can be refactored in an additive noise model, the guarantees of identifiability still hold (for the formal statement and proof see Appendix E).
@@
As we discussed, the post-ANM is generally identifiable, which suggests that the setting of Example 2 is rather artificial. This result provides the theoretical ground for training causal discovery algorithms on datasets generated from multiple identifiable SCMs. This is particularly appealing in the case of CSIvA, given the poor OOD generalization ability observed in our experiments of Section 4.2.
@@ 4.5 Can we train CSIvA on multiple causal models for better generalization?
Next, we analyze the test performance of three CSIvA networks optimized on samples from structural causal models that have different distributions for their noise terms, while keeping the mechanism types fixed. Figure 5 shows that training over different noises (beta, gamma, gumbel, exponential, mlp, uniform) always results in a network that is agnostic with respect to the noise distributions of the SCM generating the test samples, always achieving SHD $<0.1$, with the exception of datasets with mlp error terms ($0.2$ average SHD on linear and nonlinear data).
@@
We have shown that learning on mixtures of SCMs with different noise term distributions and mechanism types leads to models generalizing to a much broader class of structural causal models during testing. Hence, combining datasets generated from multiple models looks like a promising framework to overcome the limited out-of-distribution generalization abilities of CSIvA observed in Section 4.2. However, it is easier to incorporate prior assumptions on the class of causal mechanisms (linear, non-linear, post-non-linear) compared to the noise distributions (which are potentially infinite). This introduces a trade-off between amortized inference and classical methods for causal discovery: for example, RESIT, NoGAM, and CAM (Peters et al., 2014, Montagna et al., 2023c, Bühlmann et al., 2014) algorithms require no assumptions on the noise type, but only work for a limited class of mechanisms (nonlinear).
@@ 5 Conclusion
In this work, we investigate the interplay between identifiability theory and supervised learning for amortized inference of causal graphs, using CSIvA as the ground of our study. Consistent with classical algorithms, we demonstrate that good performance can be achieved if (i) we have a good prior on the structural causal model generating the test data (ii) the setting is identifiable. In particular, prior knowledge of the test distribution is encoded in the training data in the form of constraints on the structural causal model underlying their generation. With these results, we highlight the need for identifiability theory in modern learning-based approaches to causality, while past works have mostly disregarded this connection. Further, our findings provide the theoretical ground for training on observations sampled from multiple classes of identifiable SCMs, a strategy that improves test generalization to a broad class of causal models. Finally, we highlight an interesting new trade-off regarding identifiability: traditional methods like LiNGAM, RESIT, and PNL require strong restrictions on the structural mechanisms underlying the data generation (linear, nonlinear additive, or post-nonlinear) while generally being agnostic relative to the noise terms distribution. Training on mixtures of causal models instead offers an alternative that is less reliant on assumptions on the mechanisms, while incorporating knowledge about all possible noise distributions in the training data is practically impossible to achieve. We leave it to future work to reproduce our analysis on a wider class of architectures, as well as extending our study to interventional data with more than two nodes.